# Random Lévy Looptrees and Lévy Maps

### v2

@@ 1.2.   Main results
We next discuss the convergence of these continuum labelled looptrees and maps when the driving Lévy excursion varies. We will in particular identify natural families of random compact metric spaces which interpolate between the Brownian tree and the Brownian sphere. To keep this introduction as short as possible, we state here our results in the particular case where the driving Lévy process is a spectrally positive Lévy process with a drift. As was already mentioned, this case naturally appears when considering large discrete stable random planar maps, conditioned on having a fixed number of edges, vertices, and faces at the same time in a fine-tuned regime (see Section 6 for a precise statement).
Fix $\alpha\in(1,2)$ and $\vartheta\in(-\infty,\infty)$ and let $X^{\alpha,\vartheta,\mathrm{ex}}$ denote an excursion with duration $1$ of the $\alpha$-stable Lévy process with drift $-\vartheta$, characterised by the Laplace exponent $\psi(\lambda)=\lambda^{\alpha}-\vartheta\lambda$. We choose to let $-\vartheta$ denote the drift instead of $\vartheta$ in order to see the bridge version as describing the fluctuations of the stable process without drift and conditioned to end at position $\vartheta$ at time $1$. Further, let $\mathscr{L}(X^{\alpha,\vartheta,\mathrm{ex}})$ and $M^{\alpha,\vartheta}$ denote respectively the looptree coded by $X^{\alpha,\vartheta,\mathrm{ex}}$ and the corresponding Lévy map. By application of Theorem 1.1 and Theorem 1.2, all the four dimensions are equal in this case, to $\alpha$ and $2\alpha$ respectively.
The next result describes the behaviour of $X^{\alpha,\vartheta,\mathrm{ex}}$, $\mathscr{L}(X^{\alpha,\vartheta,\mathrm{ex}})$, and $M^{\alpha,\vartheta}$ for any fixed $\alpha\in(1,2)$ when ϑ±italic-$\vartheta\to\pm\infty$. It shows in particular that appropriately rescaled, the map $M^{\alpha,\vartheta}$ interpolates between the Brownian sphere and the Brownian tree as $\vartheta$ varies from $\infty$- ∞ to $\infty$. See Figure 4 for simulations of $X^{\alpha,\vartheta,\mathrm{ex}}$, and Figure 1 for simulations of $\mathscr{L}(X^{\alpha,\vartheta,\mathrm{ex}})$ and of $M^{\alpha,\vartheta}$.
Theorem 1.3.
Fix $\alpha\in(1,2)$. The following convergences in distribution hold for the Skorokhod $J_{1}$ topology for paths and the Gromov–Hausdorff–Prokhorov topology for measured metric spaces.
  1. (i)
    As $\vartheta\to-\infty$, we have:
    $s_{\alpha,\vartheta}^{1/2}\cdot X^{\alpha,\vartheta,\mathrm{ex}}\enskip\mathop{\longrightarrow}^{(d)}_{\vartheta\to-\infty}\enskip\boldsymbol{e}\qquad\text{with}\qquad s_{\alpha,\vartheta}=\frac{|\vartheta|^{(2-\alpha)/(\alpha-1)}}{(\alpha-1)\alpha^{1/(\alpha-1)}},$
    where $\boldsymbol{e}$ is the standard Brownian excursion, and further:
    $(4s_{\alpha,\vartheta})^{1/2}\cdot\mathscr{L}(X^{\alpha,\vartheta,\mathrm{ex}})\enskip\mathop{\longrightarrow}^{(d)}_{\vartheta\to-\infty}\enskip\mathscr{T}\qquad\text{and}\qquad(9s_{\alpha,\vartheta})^{1/4}\cdot M^{\alpha,\vartheta}\enskip\mathop{\longrightarrow}^{(d)}_{\vartheta\to-\infty}\enskip\mathscr{S},$
    where $\mathscr{T}$ is the standard Brownian tree coded by $\boldsymbol{e}$ and $\mathscr{S}$ the standard Brownian sphere.
  2. (ii)
    As $\vartheta\to\infty$, we have:
[…]
    as well as
    $\vartheta^{-1}\cdot\mathscr{L}(X^{\alpha,\vartheta,\mathrm{ex}})\enskip\mathop{\longrightarrow}^{(d)}_{\vartheta\to\infty}\enskip\mathscr{C}\qquad\text{and}\qquad\vartheta^{-1/2}\cdot M^{\alpha,\vartheta}\enskip\mathop{\longrightarrow}^{(d)}_{\vartheta\to\infty}\enskip\mathscr{T},$
    where $\mathscr{C}$ is the circle with unit perimeter and $\mathscr{T}$ is again the standard Brownian tree.
We interpret the convergence of the path in the second regime as $\vartheta^{-1}X^{\alpha,\vartheta,\mathrm{ex}}_{t}\to(1-t)\mathbf{1}_{\{0<t\leq 1\}}$; the latter is not right-continuous at 00, which is the reason why we reverse time in the statement.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: Simulations of excursions of $\alpha$-stable processes with a drift $-\vartheta$ with a fixed duration. Top: $(\alpha,\vartheta)=(1.5,-8)$ and $(\alpha,\vartheta)=(1.5,0)$. Bottom: $(\alpha,\vartheta)=(1.5,10)$ and $(\alpha,\vartheta)=(1.9,20)$.
Let us give some intuition concerning the asymptotic behaviour of $X^{\alpha,\vartheta,\mathrm{ex}}$. When $\vartheta\to-\infty$, any positive jump for the stable process conditioned to end at $\vartheta$ becomes very costly, and in the limit it does not jump at all, so one must expand the fluctuations (and indeed $s_{\alpha,\vartheta}\to\infty)$, which in the limit are naturally described by a Brownian bridge, hence a Brownian excursion after applying the Vervaat transform. On the other hand, the regime when $\vartheta\to\infty$ exhibits a new instance of the so-called ‘one big jump’ phenomenon [DDS08, AL11, BBY23, KR02, DJ12], in that the easiest way for the stable process to end at a large $\vartheta$ is to make a single jump of size approximately $\vartheta$, and behave microscopically compared to $\vartheta$ otherwise. It results in a so-called ‘condensation phenomenon’ already observed in several models of (loop)trees and maps [JS15, Kor15, KR19, Mar22a].
Let us note that when $\vartheta=0$, the behaviour of stable excursions and stable looptrees as $\alpha\to 1$ and $\alpha\to 2$ have been obtained in [CK14]. Our general results actually allow $\alpha$ and $\vartheta$ to vary at the same time, see precisely Theorem 6.2 and Corollary 6.3 below.
@@
In a general setting, the convergence of looptrees and maps when the driving Lévy excursion varies is discussed in Theorem 5.2 and Theorem 5.3. It is essentially established that if the transition densities of a sequence of Lévy processes converge uniformly to those of another Lévy process, then the associated labelled looptrees converge in law, and further the corresponding Lévy maps converge in law after extracting a subsequence.
Let us say a few words on the techniques used to derive these general results. First, tightness of the looptree distances is easy but that of the labels is not: we argue by adapting the proof of the usual Kolmogorov criterion and using some general concentration bounds for Gaussian processes. We then turn to the finite-dimensional marginals, that is the pairwise distances and joint labels of finitely many points in the looptree. The contribution of the large cycles can easily be controlled, and the contribution of the microscopic parts is proved by a sort of weak law of large numbers in the spirit of [CK14]. Indeed, informally, the Lévy excursion codes the length of paths in the looptree when one always follows the cycles on their right, whereas the true distance minimises left and right lengths. We show that this results in dividing the lengths by two. We actually argue on the unconditioned process first and then use a local absolute continuity argument for the excursion; the convergence of the densities is used for this transfer. Once the convergence of labelled looptrees is proved, the convergence of maps along a subsequence, and the identification of the Brownian sphere and tree in the extreme regimes, is routine as the usual arguments to control discrete maps can be adapted.
Finally, the proof of Theorem 1.3, and its actual more general version in Theorem 6.2, uses very different arguments for the Brownian excursion and the one-big jump regime. In the former case we prove the convergence of the transition densities and use the preceding paragraph. A key idea is to use an exponential change of measure to make the imposed drift $\vartheta$ ‘typical’ for the stable Lévy process. The one-big jump regime is treated very differently, and also very differently from [CK14]. We consider a large threshold, but small compared to $\vartheta$, and we show that the trajectory obtained by removing all the jumps larger than this threshold, under the bridge law and rescaled by $\vartheta^{-1}$, converges to the linear drift $-t$. This is done via Hoeffding-type concentration bounds available since the jumps are bounded. Then we show that the bridge can only make at most two jumps larger than our threshold with high probability, and then actually only one by an exchangeability argument. This shows that our rescaled bridge converges to $(\mathbf{1}_{\{U\leq t\}}-t)_{t\in[0,1]}$ where $U$ has the uniform distribution on $[0,1]$; the claim follows from the Vervaat transform. Again here, we argue first on the unconditioned process and then transfer the results to the bridge by absolute continuity. This requires to understand the asymptotic behaviour of the stable density, and we provide some new results about it, as $\alpha\to 1$, in Section 6.5.
@@ 1.3.   Open questions and plan
This work leaves open several natural questions. First, here the Lévy map $M$ is only defined as a subsequential limit of models of discrete maps, associated with discrete excursions which converge to $X^{\mathrm{ex}}$. This was the case of the Brownian sphere until uniqueness was finally proved [LG13, Mie13] and this is still the case for stable maps, for which uniqueness is under current investigation [CMR]. We believe that several geometric ideas of [CMR] can apply to prove uniqueness in the more general Lévy setting when the Gaussian parameter $\beta$ vanishes, although the technical inputs are more involved without self-similarity. When the Gaussian parameter is nonzero, it is likely that other ideas are needed.
@@
Following the first point, the convergence of looptrees and maps relies on that of the discrete excursions to $X^{\mathrm{ex}}$. The point of this paper was to build the limiting objects, which for maps requires to pass to the limit of discrete planar maps. Given a Lévy excursion, one can always make up discrete paths that converges to it and use them to build maps, which we do here. The question of proving convergence of natural discrete models of conditioned discrete excursions under optimal assumptions is left for future work.
@@
In another direction, after computing the fractal dimensions of Lévy trees [DLG05], Duquesne and Le Gall [DLG06] investigated the existence and the computation of a gauge function associated with a nontrivial Hausdorff or packing measure for the Brownian tree. Duquesne [Duq12] expressed more generally the exact packing measure of Lévy trees, and proved on the contrary that there is no (regular) Hausdorff measure, even for the stable trees [Duq10]. More recently, Le Gall [LG22] expressed the Hausdorff measure of the Brownian sphere. It is then natural to investigate these questions on the Lévy looptrees and maps.
In a last direction, in Section 6 we consider the asymptotic behaviour of the models related to $\alpha$-stable processes with a drift $-\vartheta$ as the two parameters vary. Our results cover all regimes except the two limits $(\alpha,\vartheta)\to(1,-1)$ and $(\alpha,\vartheta)\to(2,\infty)$, for which nothing can be said in general and the behaviour depends more precisely on the speed of each convergence. It is however interesting to look further at these two points, which may provide other interpolations between the two extreme behaviours displayed in Theorem 1.3.
@@
The rest of this paper is organised as follow. In Section 2 we first define formally discrete and continuous looptrees and maps. We stress again that Lévy looptrees are defined directly from the excursion $X^{\mathrm{ex}}$, whereas Lévy maps are defined as subsequential limits of discrete maps, in Theorem 2.4. Then in Section 3 we state and prove technical bounds on the volume of small balls in looptrees and maps, which provide lower bounds for the fractal dimensions. In Section 4 we first state and prove Hölder continuity estimates for the looptree and map distances, which provide upper bounds for the fractal dimensions. We then prove Theorem 1.1 and Theorem 1.2. Then Section 5 discusses the convergence of Lévy looptrees and maps when the associated processes converge. Finally Section 6 applies these results to the model of stable processes with a drift; there we state and prove extensions of Theorem 1.3.
@@ 2. Planar maps and labelled looptrees
We first define in Section 2.1 the discrete labelled looptrees and plane maps as well as their coding by discrete paths. Then we construct analogously labelled looptrees associated with (excursions of) Lévy processes in Section 2.2. We shall stay brief and refer to [Mar22b] and references therein for details. See also [Kha22] for more results on looptrees coded by a function. Finally in Section 2.3 we state and prove invariance principles, showing that the discrete objects converge to the continuum ones. We define at this occasion the Lévy maps as subsequential limits of discrete maps.
@@ 2.1.   The discrete setting
One can check that $W^{n}$ is a Łukasiewicz path with length $n+1$, i.e. a path whose increments all belong to $\mathbb{Z}_{\geq-1}=\{-1,0,1,2,\dots\}$, started from $W^{n}_{0}=0$, with $W^{n}_{n+1}=-1$ and $W^{n}_{k}\geq 0$ for every $1\leq k\leq n$. We extend it to a càdlàg path defined on the real interval $[0,n+1]$ by letting it drift at speed $-1$ between two integer times. We refer to Figure 5 for an example. This construction is invertible, we shall describe below the converse construction directly in the continuum setting. We denote by $\mathsf{L}(W^{n})$ the looptree coded by the path $W^{n}$. Let us mention that there is a well-known bijection between Łukasiewicz paths and plane trees, in which the increments of $W^{n}$ code the offspring numbers minus $1$ when performing a depth first search exploration of the tree; the tree associated with $W^{n}$ is the dotted tree on the left of Figure 2.
Refer to captionRefer to caption
@@
Let us further equip a looptree with a good labelling on the vertices. For every edge, oriented to keep the boundary on its left, we require that the difference of labels between the tip and the origin is an integer larger than or equal to $-1$. This fixes uniquely the labels of the vertices up to a global additive constant, which we prescribe by requiring the origin of the root edge to have label 00. One can again encode the labels into a discrete path $Z^{n}=(Z^{n}_{0},\dots,Z^{n}_{n})$ by following the contour of the looptree. We further extend this path continuously by linear interpolation. See again Figure 5 for an example. One can notice that the labels of the vertices of any cycle of the looptree, when read clockwise, form a bridge with increments larger than or equal to $-1$; if the cycle has length $\ell\geq 1$, there are $\binom{2\ell-1}{\ell}$ such bridges.
The bijection from [BDFG04, Section 2] once reformulated in terms of looptrees shows that looptrees equipped with a good labelling are in one-to-two correspondence with bipartite plane maps equipped with a distinguished vertex, called pointed maps. Moreover the bijection can be constructed by a simple algorithm in both directions (in addition to [BDFG04], let us refer to [Mar22b, Section 2.2]) and enjoys the following properties:
@@
Let us refer to Figure 3 for a graphical representation of this bijection. The reader acquainted with the Schaeffer bijection between labelled trees and quadrangulations can observe that the latter is a particular case of the present bijection, when every cycle of the looptree has length $2$, so the looptree is really just a tree in which every edge is doubled. In words, the construction from a labelled looptree to a map works as follows. First, shift all the labels so the minimum is $1$. Then assign to every corner of the outer face a successor, which is the first one which carries a smaller label when following the contour of the looptree, in the sense defined at the beginning of this section but extended by periodicity. The fact that the labelling is good implies that the difference of label between a corner and its successor is always exactly $-1$. The successor of the corners which carry the label $1$ is an extra vertex labelled 00 in the outer face. Then the map is obtained by removing the edges of the looptree and instead linking every corner to its successor.
@@
Let us briefly recall the model of Boltzmann maps introduced in Section 1.1. Given a sequence $(q_{k})_{k\geq 1}$, one assigns to any finite bipartite pointed map a weight given by:
$\prod_{f\text{ face}}q_{\deg(f)/2},$
where $\deg(f)$ is the degree of the face $f$. One can then sample a pointed map with $n$ edges proportionally to its weight (assuming that the total weight of $n$-edged maps, which is a finite set, is non-zero). By the previous bijection, this transfers to sampling a looptree with $n$ edges and equipped with a good labelling proportionally to the weight $\prod_{c\text{ cycle}}q_{\text{length}(c)}$. Recall that each cycle $c$ with length $\ell\geq 1$ in a looptree can be labelled in $\binom{2\ell-1}{\ell}$ different ways. Then the looptree, without the labels, is sampled proportionally to the weight:
@@
and then conditionally on it, the labels on each cycle are uniform random bridges with increments larger than or equal to $-1$, independent of each other. Recall that the length of the cycles of the looptree equal the size plus one of the nonnegative increments of the associated Łukasiewicz path. This Łukasiewicz path, with duration $n+1$, is thus sampled proportionally to the weight:
@@
Let us mention that this paper focuses on the continuum objects, so we consider pointed maps with $n$ edges which is the simplest model to study. However one may consider maps with $n$ vertices or $n$ faces instead, or even more general notions of size, see e.g. [Mar22a, Section 6.4] and references therein, or combine them and condition the map by its number of vertices, edges, and faces at the same time [KM23]. In another direction, one can consider maps without any distinguished vertex. Note that when we do not condition on the number of vertices, distinguishing a vertex in the map leads to a size-biasing of the latter. However standard arguments in the theory allow to show that the effect of this size-biasing disappears in the limit, provided some technical inputs, see e.g. [Mar22a, Proposition 6.4] and references therein. Finally one can consider maps with a boundary [Bet15, BM17] by prescribing the degree of the face to the right of the root edge in the map, and hence the length of the cycle to the right of the root edge in the looptree, so finally the size of the first increment of the Łukasiewicz path. By removing this first jump, one obtains a so-called first-passage bridge, and all this present work extends readily to this setting, see [Mar22a, Mar22b]. We simply consider maps without boundary to ease the notation.

### v3

@@ 1.2.   Main results
In a general setting, the convergence of looptrees and maps when the driving Lévy excursion varies is discussed in Theorem 6.2 and Theorem 6.3. It is essentially established that if the transition densities of a sequence of Lévy processes converge uniformly to those of another Lévy process, then the associated labelled looptrees converge in law, and further the corresponding Lévy maps converge in law after extracting a subsequence. A typical example of application we have in mind is the case where $X^{(\lambda)}$ is an $\alpha$-stable spectrally positive Lévy process with drift $\lambda$, with varying $\alpha$ and $\lambda$. This will be studied in a companion paper.
@@ 1.3.   Open questions and plan
This work leaves open several natural questions. First, here the Lévy map $M$ is only defined as a subsequential limit of models of discrete maps, associated with discrete excursions that converge to $X^{\mathrm{ex}}$. This was the case for the Brownian sphere in [LG07] until uniqueness was finally proved [LG13, Mie13], and similarly for stable maps [LGM11], for which uniqueness has been proved very recently [CMR25]. We believe that several geometric ideas of [CMR25] can apply to prove uniqueness in the more general Lévy setting when the Gaussian parameter $\beta$ vanishes, although the technical inputs are more involved without self-similarity. When the Gaussian parameter is nonzero, it is likely that other ideas are needed.
@@
Following the first point, the convergence of looptrees and maps relies on that of the discrete excursions to $X^{\mathrm{ex}}$. The point of this paper was to construct the limiting objects related to $X^{\mathrm{ex}}$. Given a Lévy excursion, one can always make up discrete paths that converge to it and use them to build discrete maps, which we do here. The question of proving convergence of natural discrete models of maps and conditioned discrete excursions under optimal assumptions is left for future work.
@@
In another direction, after computing the fractal dimensions of Lévy trees [DLG05], Duquesne and Le Gall [DLG06] investigated the existence and the computation of a gauge function associated with a nontrivial Hausdorff or packing measure for the Brownian tree. Duquesne [Duq12] expressed more generally the exact packing measure of Lévy trees, and proved on the contrary that there is no (regular) Hausdorff measure, even for the stable trees [Duq10]. More recently, Le Gall [LG22] expressed the Hausdorff measure of the Brownian sphere. We plan to investigate these questions on the Lévy looptrees and maps.
@@
The rest of this paper is organised as follow. In Section 2 we first define formally discrete and continuous looptrees and maps. We stress again that Lévy looptrees are defined directly from the excursion $X^{\mathrm{ex}}$, whereas Lévy maps are defined as subsequential limits of discrete maps, in Theorem 2.3. Section 3 applies these results to the model of stable processes with a drift. Then in Section 4 we state and prove technical bounds on the volume of small balls in looptrees and maps, which provide lower bounds for the fractal dimensions. In Section 5 we first state and prove Hölder continuity estimates for the looptree and map distances, which provide upper bounds for the fractal dimensions. We then prove Theorem 1.1 and Theorem 1.2. Finally, Section 6 discusses the convergence of Lévy looptrees and maps when the associated processes converge.
Acknowledgment
The authors would like to thank the associated editor and the referees for their detailed comments and suggestions that greatly improved this paper.
@@ 2. Planar maps and labelled looptrees
We first define in Section 2.1 the discrete labelled looptrees and plane maps as well as their coding by discrete paths. Then we construct analogously labelled looptrees associated with (excursions of) Lévy processes in Section 2.2. We shall stay brief and refer to [Mar24] and references therein for details. See also [Kha22] for more results on looptrees coded by a function. Finally in Section 2.3 we state and prove invariance principles, showing that the discrete objects converge to the continuum ones. We define at this occasion the Lévy maps as subsequential limits of discrete maps.
@@ 2.1.   The discrete setting
One can check that $W^{n}$ is a Łukasiewicz path with length $n+1$, i.e. a path whose increments all belong to $\mathbb{Z}_{\geq-1}=\{-1,0,1,2,\dots\}$, started from $W^{n}_{0}=0$, with $W^{n}_{n+1}=-1$ and $W^{n}_{k}\geq 0$ for every $1\leq k\leq n$. We extend it to a càdlàg path defined on the real interval $[0,n+1]$ by letting it drift at speed $-1$ between two integer times. We refer to Figure 4 for an example. This construction is invertible, we shall describe below the converse construction directly in the continuum setting. We denote by $\mathsf{L}(W^{n})$ the looptree coded by the path $W^{n}$. Let us mention that there is a well-known bijection between Łukasiewicz paths and plane trees, in which the increments of $W^{n}$ code the offspring numbers minus $1$ when performing a depth first search exploration of the tree; the tree associated with $W^{n}$ is the dotted tree on the left of Figure 2. It is important to note that $\mathsf{L}(W^{n})$ is not the looptree naturally associated with the plane tree coded by $W^{n}$ as considered in [CK14, Kha22] (each internal vertex of the tree with its right-most offspring should be merged to obtain $\mathsf{L}(W^{n})$). This version however is the one primarily studied in [Mar24] and that appeared previously in [CK15, KR20, Ric18] under the notation $\overline{\mathsf{Loop}}$ in view of applications to random planar maps. This seemingly small difference actually affects the continuous part in the scaling limits, that is, the constant $a$ in (2.6) below.
@@
When considering scaling limits of looptrees, it will be useful to have a canonical order of their vertices. First for each $0\leq k\leq n$, let $u^{n}_{k}$ denote the origin of the oriented edge $e_{k}$ in the contour sequence. This list contains redundancies that we can remove by only keeping the last appearance of each vertex to extract a list of the vertices $(v^{n}_{1},\dots,v^{n}_{N})$ where $N$ is the total number of vertices; in particular, the root vertex is $u^{n}_{0}=u^{n}_{n}=v^{n}_{N}$ is the origin of the root edge. The reason to keep the last appearance and not the first one come from the coding from the Łukasiewicz path $W^{n}$: the vertices are in one-to-one correspondence with the negative increments of the path, so $N$ equals the number of such increments, and the time of a negative jump is precisely the last appearance of the corresponding vertex in the list $(u^{n}_{0},\dots,u^{n}_{n})$.
@@
Let us further equip a looptree with a good labelling on the vertices. For every edge, oriented to keep the boundary on its left, we require that the difference of labels between the tip and the origin is an integer larger than or equal to $-1$. These differences, together with the label of the origin of the root edge fix uniquely the labels of all the vertices. One can again encode the labels into a discrete path $Z^{n}=(Z^{n}_{0},\dots,Z^{n}_{n})$ by following the contour of the looptree. We further extend this path continuously by linear interpolation. See again Figure 4 for an example. One can notice that the labels of the vertices of any cycle of the looptree, when read clockwise, form a bridge with increments larger than or equal to $-1$; if the cycle has length $\ell\geq 1$, there are $\binom{2\ell-1}{\ell}$ such bridges.
The bijection from [BDFG04, Section 2] once reformulated in terms of looptrees shows that looptrees equipped with a good labelling are in one-to-two correspondence with bipartite plane maps equipped with a distinguished vertex, called pointed maps. Moreover the bijection can be constructed by a simple algorithm in both directions (in addition to [BDFG04], let us refer to [Mar24, Section 2.2]) and enjoys the following properties:
@@
Let us refer to Figure 3 for a graphical representation of this bijection. The reader acquainted with the Schaeffer bijection between labelled trees and quadrangulations can observe that the latter is a particular case of the present bijection, when every cycle of the looptree has length $2$, so the looptree is really just a tree in which every edge is doubled.
In short, the construction of a map from a labelled looptree works as follows. First, shift all the labels so the minimum is $1$. Then assign to every corner of the outer face a successor, which is the first one which carries a smaller label when following the contour of the looptree, in the sense defined at the beginning of this section but extended by periodicity. The fact that the labelling is good implies that the difference of label between a corner and its successor is always exactly $-1$. The successor of the corners which carry the label $1$ is an extra vertex labelled 00 in the outer face. Then the map is obtained by removing the edges of the looptree and instead linking every corner to its successor. The extra vertex 00 is the distinguished vertex of the map, and the labels of the vertices indeed correspond to their distance in the map to this vertex: the chain of successors forms a geodesic path to 00.
This construction allows to canonically order the vertices of a map. Recall the order $(v^{n}_{1},\dots,v^{n}_{N})$ of the vertices of a looptree with $n$ edges and $N$ vertices. Then these are all the vertices of the associated map except the distinguished one which we set as $v^{n}_{0}$. In addition, the vertex $v^{n}_{N}$ corresponds to the origin of the root edge in the looptree; in the map, it is an extremity of the root edge, but not necessarily its origin, but rather the one farther away from $v_{0}^{n}$.
@@
Let us briefly recall the model of Boltzmann maps introduced in Section 1.1. Let $\deg(f)$ is the degree of the face $f$, then given a sequence $(q_{k})_{k\geq 1}$, one assigns to any finite bipartite pointed map a weight given by:
$\prod_{f\text{ face}}q_{\deg(f)/2}.$
One can then sample a pointed map with $n$ edges proportionally to its weight (assuming that the total weight of the finite set of $n$-edged maps is non-zero). By the previous bijection, this transfers to sampling a looptree with $n$ edges and equipped with a good labelling proportionally to the weight $\prod_{c\text{ cycle}}q_{\text{length}(c)}$. Recall that each cycle $c$ with length $\ell\geq 1$ in a looptree can be labelled in $\binom{2\ell-1}{\ell}$ different ways. Then the looptree, without the labels, is sampled proportionally to the weight:
@@
and then conditionally given this looptree, the labels on each cycle are uniform random bridges with increments larger than or equal to $-1$, independent of each other. Recall that the length of the cycles of the looptree equal the size plus one of the nonnegative increments of the associated Łukasiewicz path. This Łukasiewicz path, with duration $n+1$, is thus sampled proportionally to the weight:
@@
Let us mention that this paper focuses on the continuum objects, so we consider pointed maps with $n$ edges which is the simplest model to study. However one may consider maps with $n$ vertices or $n$ faces instead, or even more general notions of size, see e.g. [Mar22, Section 6.4] and references therein, or combine them and condition the map by its number of vertices, edges, and faces at the same time [KM23]. In another direction, one can consider maps without any distinguished vertex. Note that when we do not condition on the number of vertices, distinguishing a vertex in the map leads to a size-biasing of the latter. However standard arguments in the theory allow to show that the effect of this size-biasing disappears in the limit, provided some technical inputs, see e.g. [Mar22, Proposition 6.4] and references therein. Finally one can consider maps with a boundary [Bet15, BM17] by prescribing the degree of the face to the right of the root edge in the map, and hence the length of the cycle to the right of the root edge in the looptree, so finally the size of the first increment of the Łukasiewicz path. By removing this first jump, one obtains a so-called first-passage bridge, and all this present work extends readily to this setting, see [Mar22, Mar24]. We simply consider maps without boundary to ease the notation.